OpenStudy (anonymous):

(3x+1)=(2x+2) it's asking for work, but I don't know how. Please help

OpenStudy (anonymous):

(3x+1)=(2x+2) 3x= 2x+1 x=1

OpenStudy (mathstudent55):

I assume you need to solve the equation for x. First step is to drop the parentheses. They are not needed. 3x + 1 = 2x + 2

OpenStudy (anonymous):

Thanks, but how do I get to it?

OpenStudy (anonymous):

Alright

OpenStudy (mathstudent55):

Now you want all terms with x on the left side, so we subtract 2x from both sides.

OpenStudy (anonymous):

Alright 3x+1=2x+2 1X+1=x+2

OpenStudy (mathstudent55):

3x + 1 = 2x + 2 -2x -2x ---------------- x + 1 = 2 Now we have x + 1 = 2 We want x alone on the left side, so we subtract 1 from both sides: x + 1 = 2 - 1 -1 ------------ x = 1 The answer is x = 1

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (mathstudent55):

No, your last response is not correct. When you subtract 2x from both sides, there is no x left on the right side. 3x + 1 = 2x+2 -2x -2x ------------------- x + 1 = 2

OpenStudy (anonymous):

Wait, what type of equation is this?

OpenStudy (anonymous):

Thank you for that little explanation. I haven't got into this type of math in a while and I don't remember some of the basics @mathstudent55

OpenStudy (mathstudent55):

Now that we have x = 1, we can check to see if the solution is correct. Take the original equation, (3x + 1) = (2x + 2), and replace x with our solution, 1. 3(1) + 1 = 2(1) + 2 3 + 1 = 2 + 2 4 = 4 Since 4 = 4 is a true statement, that means our solution, x = 1, is the correct solution.

OpenStudy (mathstudent55):

You're welcome.